OpenStudy (anonymous):

-4<2x-6 less than or equal to 8

OpenStudy (lgbasallote):

\[-4<2x - 6 \le 8\] add 6 to all sides \[\implies -4 + 6 < 2x - 6 + 6 \le 8 + 6\] combine like terms \[\implies 2 < 2x \le 14\] divide all sides by 2 \[\implies \frac 22 < \frac{2x}{2} \le \frac{14}{2}\] simplify \[\implies 1 < x \le 7\] does that help?